# Emergence of Chimeras States in One-dimensional Ising model with Long-Range Diffusion

### v1

@@ 3.2.3 Phase diagram of the system
The analysis of both phase diagrams shows the emergence of three main regions: Region A which corresponds to $\ell_{dom}$ $\approx 0.5$, i.e., wide attractors, with very short $\tau_{ss}$, corresponding to two alternating domains (positive and negative) that form rapidly. Secondly, Region B where $\ell_{dom}$ declines gradually and $\tau_{ss}$ increases gradually with $r$, indicating the presence metastable chimera domains coexisting with attractors. Finally, Region C is characterized by an abrupt drop in $\ell_{dom}$ and abrupt increase in $\tau_{ss}$, matching the $|m_{0}|=2r-1$ criterion from Eq. 3.7, as represented by the dashed red line. In Region C, attractors can not exist so only chimeras remains. These numerical results confirm the analytical predictions (see also Eq. 3.7) for attractor formation, as well showing the metastable region discussed in Section 3.2.2.
@@
The long‐lived metastable chimera states in Region B can be understood by mapping the noisy domain of $n^{+}$ positive spins onto a constrained random walk of $n^{+}$ walkers within the overlapping “ball” $B_{n^{+}}$ (see definition 3.2). At zero‐temperature, each positive spin executes exchanges with an available negative neighbor in $B_{n^{+}}$ with equal probability (see Eq. 3.4). As $n^{+}/R$ increases ($|m_{0}|$ decreases) for a fixed $r$, $|B_{n^{+}}|$ shrinks in size relative to $n^{+}$ and one arrives to a more “crowded” random‐walk habitat: walkers (positive spins) must coordinate over fewer available sites, and their joint diffusion slows down. Chimeras would then take longer to find another chimera (forming a domain that, in many cases, is larger than $R+1$ and then would decay into an attractor instead) or find an attractor and being absorbed into it. The behavior in Region B can be described in terms of a mesoscopic scale: the system would first organize into moving chimeras and fixed attractors (depending on the number of positive spins that are in a certain medium size or mesoscopic region) and then these chimeras would move, interacting with other moving chimeras or some attractors.
@@ 4 Conclusions
An essential factor in the emergence of chimera states is that the Metropolis transition rate (Eq. 2.4) inherently accepts zero-energy moves. In classical systems, however, any transition involves a finite energy cost, raising the question of why such moves are accepted. The simulated dynamics should therefore not be interpreted as physically realizable. Rather, unconditional acceptance of transitions between configurations with identical energy (and thus identical Boltzmann weights, $e^{-\beta\mathcal{H}}$) ensures unbiased sampling of the configuration space. Importantly, the main result – the existence of stable chimera states in the one-dimensional Ising model with long-range interactions – remains robust when other transition rates, such as the Glauber dynamics, are employed (data not shown).
In conclusion, we have shown that a uniformly coupled, one-dimensional Ising system supports the existence of chimera states under diffusive dynamics in the thermodynamic limit. An interesting question to address is if such patterns can also appear and be stable in non-equilibrium situations, for example including reaction and diffusion competing dynamics. Extending the present study for such situations could be of great interest. In fact, some preliminary simulations of such dynamics point out in this direction (data not shown). However, we think that a fully study of this non-equilibrium system setup is beyond the scope of the present study and be more appropriate for a future work.

### v2

@@ 3.2.3 Phase diagram of the system
The analysis of both phase diagrams shows the emergence of three main regions: Region A which corresponds to $\ell_{dom}$ $\approx 0.5$, i.e., wide attractors, with very short $\tau_{ss}$, corresponding to two alternating domains (positive and negative) that form rapidly. Secondly, Region B where $\ell_{dom}$ declines gradually and $\tau_{ss}$ increases gradually with $r$, indicating the presence metastable chimera domains coexisting with attractors. Finally, Region C is characterized by an abrupt drop in $\ell_{dom}$ and abrupt increase in $\tau_{ss}$, matching the $|m_{0}|=2r-1$ criterion from Eq. (12), as represented by the dashed red line. In Region C, attractors can not exist so only chimeras remains. These numerical results confirm the analytical predictions (see also Eq. (12)) for attractor formation, as well showing the metastable region discussed in Section 3.2.2.
@@
The long‐lived metastable chimera states in Region B can be understood by mapping the noisy domain of $n^{+}$ positive spins onto a constrained random walk of $n^{+}$ walkers within the overlapping “ball” $B_{n^{+}}$ (see definition (7)). At zero‐temperature, each positive spin executes exchanges with an available negative neighbor in $B_{n^{+}}$ with equal probability (see Eq. (9)). As $n^{+}/R$ increases ($|m_{0}|$ decreases) for a fixed $r$, $|B_{n^{+}}|$ shrinks in size relative to $n^{+}$ and one arrives to a more “crowded” random‐walk habitat: walkers (positive spins) must coordinate over fewer available sites, and their joint diffusion slows down. Chimeras would then take longer to find another chimera (forming a domain that, in many cases, is larger than $R+1$ and then would decay into an attractor instead) or find an attractor and being absorbed into it. The behavior in Region B can be described in terms of a mesoscopic scale: the system would first organize into moving chimeras and fixed attractors (depending on the number of positive spins that are in a certain medium size or mesoscopic region) and then these chimeras would move, interacting with other moving chimeras or some attractors.
@@ 4 Conclusions
An essential factor in the emergence of chimera states is that the Metropolis transition rate (Eq. (4)) inherently accepts zero-energy moves. In classical systems, however, any transition involves a finite energy cost, raising the question of why such moves are accepted. The simulated dynamics should therefore not be interpreted as physically realizable. Rather, unconditional acceptance of transitions between configurations with identical energy (and thus identical Boltzmann weights, $e^{-\beta\mathcal{H}}$) ensures unbiased sampling of the configuration space. Importantly, the main result – the existence of stable chimera states in the one-dimensional Ising model with long-range interactions – remains robust when other transition rates, such as the Glauber dynamics, are employed (data not shown).
In conclusion, we have shown that a uniformly coupled, one-dimensional Ising system supports the existence of chimera states under diffusive dynamics in the thermodynamic limit. Although we have performed our study in noiseless conditions ($T=0$) where simple analytical derivations can be obtained, preliminary simulations shows that most of the behaviour reported in the present work also holds for $T>0$, including a rich phase diagram with the appearance of stable chimeras and attractors but now with boundaries between chimeras and attractors more diffuse and with the convergence of chimeras into attractors occurring quickly. However, we think that the extension of the present work for $T>0$ is more appropriate for a future work since it requires an in-depth analysis and the development of different types of analytical approaches for the proper characterization of the system’s emergent behavior. Another interesting question to address is if such intriguing chimera patterns can also appear and be stable in non-equilibrium situations, for example including reaction and diffusion competing dynamics. Extending the present study for such situations could be of great interest. In fact, some preliminary simulations of such dynamics point out in this direction (data not shown). We think, however, that a fully study of this non-equilibrium system setup is beyond the scope of the present study and be more appropriate for a future work.